OpenStudy (anonymous):

solve for x; x+12 over 4 + x-2 over 16 = x over 4

OpenStudy (anonymous):

Can you try to use parenthesis, please? It's hard to know where each function is.

OpenStudy (anonymous):

(x+12) / (4) + (x-2) / (16) = (x) over (4)

OpenStudy (anonymous):

is that what you mean.?

OpenStudy (anonymous):

((x+12)/4)+((x-2)/16)=x/4 by taking lcm (4(x+12)+(x-2))/16=x/4............. 4x+48+x-2=x*16/4 .............opening brakets nd *ing both sides by 16 5x-46=4x x-46=0 x=46

OpenStudy (anonymous):

thankyou very much :)

OpenStudy (anonymous):

lolx thnks again

OpenStudy (anonymous):

hey r u thr

OpenStudy (anonymous):

what.?

OpenStudy (anonymous):

wats ur real name:::

OpenStudy (anonymous):

why... ? lol

OpenStudy (anonymous):

for just to start conversation:P

OpenStudy (anonymous):

emily.

OpenStudy (anonymous):

nice one

OpenStudy (anonymous):

meaning?

OpenStudy (anonymous):

???????????????????????????